The consecutive sides of an equiangular hexagon have measures of 5, 4, 9, 3, 6 and n centimeters. What is the value of n?
Answer by ikleyn(52754)

.
Extend the sides of the hexagon until they meet. You should get big equilateral triangle and three small equilateral triangles.


MAKE A SKETCH !

Write the equation for the large triangle side lengths: 5 + 4 + 9 = 5 + n + 6, which implies n = 4 + 9 - 6 = 7.

Answer.   The value of  "n"  is equal to  7.